The lenght of a rectangle is 3 time the width. The perimeter is 120m. Find the lenght and width
Answer by checkley77(12844)

L=3W
2L+2W=120
2*3W+2W=120
6W+2W=120
8W=120
W=120/8
W=15 M. IS THE WIDTH.
L=3*15=45 M. FOR THE LENGTH.
pROOF:
2*45+2*15=120
90+30=120
120=120